# Robust Bayesian Scene Reconstruction by Leveraging Retrieval-Augmented Priors

### v2

@@ Abstract
Constructing 3D representations of object geometry is critical for many downstream robotics tasks, particularly tabletop manipulation problems. These representations must be built from potentially noisy partial observations. In this work, we focus on the problem of reconstructing a multi-object scene from a single RGBD image, generally from a fixed camera in the scene. Traditional scene representation methods generally cannot infer the geometry of unobserved regions of the objects from the image. Attempts have been made to leverage deep learning to train on a dataset of observed objects and representations, and then generalize to new observations. However, this can be brittle to noisy real-world observations and objects not contained in the dataset, and cannot reason about their confidence. We propose BRRP, a reconstruction method that leverages preexisting mesh datasets to build an informative prior during robust probabilistic reconstruction. In order to make our method more efficient, we introduce the concept of retrieval-augmented prior, where we retrieve relevant components of our prior distribution during inference. The prior is used to estimate the geometry of occluded portions of the in-scene objects. Our method produces a distribution over object shape that can be used for reconstruction or measuring uncertainty. We evaluate our method in both simulated scenes and in the real world. We demonstrate the robustness of our method against deep learning-only approaches while being more accurate than a method without an informative prior.
@@ I Introduction
The ability to construct internal representations of its operating environment is key for robot autonomy. These representations need to be particularly fine-grained for robotic manipulation, which often requires closely interacting with and avoiding objects. These interactions make it necessary for robots to develop an understanding of the geometry within their vicinity. Explicit 3D representations of the geometry of the scene are often required for the robust usage of downstream grasping and motion planning algorithms. These representations must be built from observations that are both noisy and, due to occlusion, only contain partial information of the scene. In our case, we focus on the problem of building a 3D representation of multi-object scenes from a single RGBD camera image.
@@
Deep learning methods for 3D reconstruction generally lack a sense of uncertainty about the shape of the object. This can be particularly detrimental when an object in the scene is only partially observed and we seek to factor in the geometry of observed and occluded regions. Such uncertainty can enable safer and more robust operation in a range of downstream tasks, such as robot grasping [1, 2], safe motion generation and active learning.
@@
Another common approach to building full 3D representations from a partial view is to build a representation, solely from the observed data without considering prior information from mesh datasets. A common example of this is the Gaussian Process Implicit Surface [3] model. A more recent example is V-PRISM [4], which probabilistically maps tabletop scenes without using prior information. These reconstruction methods are more robust to unknown objects because they do not rely on any training distribution. However, this also means that they cannot reconstruct the unobserved backside of objects. Humans have the remarkable ability to infer the geometry of scenes based on prior experience, we seek to imbue robots with the same capability.
Figure 1: Our method, BRRP, (a), (b) takes an input segmented RGBD image, and (c) retrieves objects to use as a prior, which allow it to (d) reconstruct the scene as well as (e) capture principled uncertainty about object shape.
@@
In this work, we introduce a novel Bayesian approach for robustly reconstructing multi-object tabletop scenes by leveraging object-level shape priors. We present Bayesian Reconstruction with Retrieval-augmented Priors (BRRP). BRRP is resilient to many of the pitfalls of learning-based methods while still being able to leverage an informative prior to more accurately reconstruct known objects. To further improve efficiency, we introduce the idea of a retrieval-augmented prior, where we retrieve relevant components of our prior distribution based on classification results. We begin with an observed RGBD image with corresponding instance segmentations. Then, we compute an identification result that predicts which objects from our database should be retrieved to use as a prior during reconstruction. We use this prior along with a sampled likelihood to infer a posterior distribution over object shapes. Because we solve for a distribution, we can recover principled uncertainty about each object’s shape. In practice, we use a pre-existing foundation model to perform the identification and use registration to make our prior pose invariant. An example scene and reconstruction can be seen in Figure 1.
@@
We conduct experiments on BRRP in both procedurally generated scenes and in the real world. We quantitatively show that BRRP results in accurate reconstructions and that BRRP is robust to unknown objects in the generated scenes. We qualitatively show that BRRP is robust to noisy real world scenes collected from an RGBD camera, and is able to capture uncertainty within the reconstructions.
Our three primary contributions can be summarized as follows:
@@
Our paper has the following organization. We overview related works in Section II. In Section III we cover the necessary mathematical preliminaries for our method. In Section IV, we introduce the concept of retrieval-augmented priors. Our method, BRRP, is introduced in Section V. Then, we perform experiments in Section VI before a conclusion in Section VII
@@ II Related Works
3D Representations. There are many different ways of representing 3D geometry of a scene. In the mapping literature, techniques such as truncated signed distance functions [5] build voxelized representations of an environment. Hilbert maps [6], on the other hand, are a continuous occupancy map that takes the form of a linear function over some hinge point feature space. Hilbert map representations have also been extended to Bayesian Hilbert maps of various forms [7, 8, 9]. Neural implicit functions have also been used to represent continuous 3D geometry [10, 11, 12]. Other representations are built using differentiable rendering and combining multiple views, especially for novel-view synthesis. A Neural radiance field (NeRF) [13] is an example of this, in which a neural network maps 3D position to density and color. 3D Gaussian splatting [14] tackles a similar but with a set of Gaussians instead of a neural network. Foundation models have also been developed for this task in [15] and applied to robotics [16]. Other representation primitives have been studied, including super quadrics [17].
@@
3D Reconstruction with Deep Learning. Many methods have been proposed as ways to leverage deep learning to reconstruct scenes or objects. While some methods aim to predict object shape from RGB data only [18, 19, 20, 21], we instead focus on using depth measurements during reconstruction. DeepSDF [10] is a method to reconstruct an object by running inference-time optimization to recover a latent code for a neural implicit function. In the context of robotics, [22] extends DeepSDF to have uncertainty-awareness. Other work, such as occupancy networks [11] or PointSDF [23] try to directly predict such a latent code without inference-time optimization. Deep learning has also been leveraged to learn kernels, which are used to construct a continuous signed distance function [24, 25]. Language is also used during reconstruction in [26] and [27]. Another work uses a voxel-to-voxel variational autoencoder conditioned on bounding boxes [28]. While these works typically focus on single object scenes, other work focuses on reconstruction scenes with multiple occluding objects. For example, [29] learns a reconstruction from a voxel representation with different channels to account for occlusion. Another method uses silhouettes to refine the initial reconstructions [30]. In practice, these deep learning approaches can struggle to reconstruct noisy scenes with multiple, highly occluded, unknown objects on real world depth cameras. Inaccurate segmentation can also be a problem for many of these methods as well.
@@
3D Reconstruction without Deep Learning. There are also many approaches to perform probabilistic 3D reconstruction without deep learning. Some of such methods for reconstruction use informative prior information by assuming fixed classes of objects, such as 3DP3 [31]. Other methods use an uninformative prior, such as Gaussian process implicit surfaces (GPIS) [3]. While there is an extension of GPIS to a slightly more informative prior [32], the only priors that can be enforced are specifically spherical, ellipsoidal, cylidrical, or planar priors. In our work, we derive our prior from pre-existing mesh datasets. V-PRISM [4] is another method that probabilistically maps the scene using a multi-class framing.
@@
Using Reconstructions in Manipulation. 3D reconstruction methods have seen extensive use in manipulation. In [23], PointSDF provides collision constraints during grasping. PointSDF is also utilized in [33], where tactile sensors are used along with the reconstruction during grasping. A learning-based voxel representation is used for grasping in [34]. Neural shape completion is also used during the anthropomorphic grasping pipeline proposed in [35]. GPIS is also a common representation for manipulation applications. Some recent work has utilized the uncertainty from GPIS representations during grasp selection [1, 2]. We believe BRRP provides principled uncertainty measurements that can similarly be utilized in downstream manipulation tasks.
@@ III-A Hilbert Maps
A Hilbert map [6] is a continuous occupancy map. It represents the environment by a continuous function that is defined by a linear function of a fixed feature transform. Typically this feature transform is induced by a set of hinge points, $\{\mathbf{h}_{1},...,\mathbf{h}_{H}\}\subset\mathbb{R}^{3}$ and a translation-invariant kernel $k(d)$. The transform is then defined as:
@@
Figure 2: (a) A hinge point feature transform induced by a set of hinge points is used by Hilbert maps [6]; (b) these maps are built by first sampling negative samples along the unoccupied portions of the camera ray.
@@
Figure 2 shows an illustration of both the hinge point feature transform and the negative samples used during Hilbert map construction.
Hilbert maps have previously been extended to Bayesian Hilbert maps, where a distribution over weights is modeled as a multivariate Gaussian [7, 8, 9]. There is also a multiclass variant that defines a weight matrix with Gaussian rows described in [4]. In this work, we adopt the Hilbert map representation, but model the distribution over weights as a collection of particles. This allows our method the capability to capture irregular, non-Gaussian posterior distributions.
@@ III-B Stein Variational Gradient Descent
Figure 3: Overview of BRRP method. We begin with a segmented RGBD image and (a) feed cropped images of each segment into CLIP to get object probabilities (Section V-B) Then, we retrieve and (b) register the the top-k objects in the prior. This gives us a set of registered prior samples (Section V-A). We also (c) compute negative samples based on the observed segmented point cloud (Section V-C). Finally, (d) we run SVGD optimization to recover a posterior distribution over Hilbert map weights (Section V-D). We can use this distribution to both reconstruct the scene as well as measure uncertainty.
Stein Variational Gradient Descent (SVGD) [36] is an algorithm for variational inference that closely resembles gradient descent. The general problem of variational inference is to find a distribution $q^{*}\in\mathcal{Q}$ that is close to some target distribution $p$. Usually, this takes the form of an optimization problem over the Kullback-Leibler (KL) divergence:
@@
SVGD aims to iteratively transform $q$ in descent directions of the KL divergence in a $d$-dimensional reproducing kernel Hilbert space (RKHS), $\mathcal{H}^{d}$. Because this Hilbert space is a space of functions, a descent direction requires deriving the functional gradient of our KL divergence objective.
@@
From [36]. Let $T(\mathbf{x})=\mathbf{x}+f(\mathbf{x})$, where $f\in\mathcal{H}^{d}$ and $q_{[T]}$ is the density of random variable $\mathbf{z}=T(\mathbf{x})$ when $\mathbf{x}\sim q$. Then
@@
The result of these iterations is that the set of particles converges to an approximation of the target distribution $p$. Importantly, Equation 2 only relies on the gradient of the log of $p$, which means we can perform variational inference to an unnormalized distribution. Such unnormalized distributions are commonplace in many Bayesian inference problems, including ours.
@@ IV Retrieval-Augmented Priors
Retrieval-augmented generation [37] was originally introduced in the context of improving language generation. The work has served as inspiration for an approach to affordance-prediction in [38]. In our case, we draw inspiration from retrieval-augmented generation, but we use the retrieved results to improve efficiency in certain explicit formulations for prior density functions during Bayesian inference.
To motivate retrieval-augmented priors, consider the problem of Bayesian inference with a mixture model acting as the prior distribution. Given some data, we would like to infer a posterior distribution over hypotheses. If we have a mixture model as a prior distribution, then:
$P(H|D)\propto P(D|H)\sum_{c=1}^{C}P(H|c).$ (3)
If our prior distribution has a lot of components, it may be inefficient to fully evaluate. This could be a serious problem for algorithms like SVGD, which requires iteratively computing the gradient of both the likelihood and prior. Inspired by [37], the insight behind retrieval-augmented priors is to determine which subset of the prior distribution components to retrieve and use given some detection result $R$. Conditioning on this detection result, we have a new posterior distribution, $P(H|D,R)$. Making an independence assumption,
@@
$P(H|D,R)\propto P(D|H)\sum_{c\in\text{topk}}P(H|c)P(c|R)$ (4)
This means that we only need to evaluate a subset of the prior distribution components.

### v3

# Robust Bayesian Scene Reconstruction with Retrieval-Augmented Priors for Precise Grasping and Planning

@@ Abstract
Constructing 3D representations of object geometry is critical for many robotics tasks, particularly manipulation problems. These representations must be built from potentially noisy partial observations. In this work, we focus on the problem of reconstructing a multi-object scene from a single RGBD image using a fixed camera. Traditional scene representation methods generally cannot infer the geometry of unobserved regions of the objects in the image. Attempts have been made to leverage deep learning to train on a dataset of known objects and representations, and then generalize to new observations. However, this can be brittle to noisy real-world observations and objects not contained in the dataset, and do not provide well-calibrated reconstruction confidences. We propose BRRP, a reconstruction method that leverages preexisting mesh datasets to build an informative prior during robust probabilistic reconstruction. We introduce the concept of a retrieval-augmented prior, where we retrieve relevant components of our prior distribution from a database of objects during inference. The resulting prior enables estimation of the geometry of occluded portions of the in-scene objects. Our method produces a distribution over object shape that can be used for reconstruction and measuring uncertainty. We evaluate our method in both simulated scenes and in the real world. We demonstrate the robustness of our method against deep learning-only approaches while being more accurate than a method without an informative prior. Through real-world experiments, we particularly highlight the capability of BRRP to enable successful dexterous manipulation in clutter.
@@ I Introduction
To operate autonomously in novel settings, a robot must construct rich internal representations of its environment. These representations need to be particularly fine-grained for robotic manipulation, which often requires closely interacting with and avoiding objects. Consider the task of robotic grasping in clutter; a robot needs to avoid unwanted collisions, while also operating in close proximity to one or more objects to make the necessary contact for the grasp. State of the art grasp synthesis methods often rely on motion planners to reach grasp poses in clutter [chen2024springgrasp, de2024task, matak2022planning, martin-arxiv-fpte], which need accurate 3D geometry for collision checking. Thus, a lack of geometric understanding can lead to unwanted collisions causing task failure or worse damage to the robot or environment. This is pictured in Figure 1. Many grasping and motion planning algorithms require explicit 3D representations of the scene’s geometry. These representations must be built from observations that are both noisy and, due to occlusion, only contain partial information of the scene. In our case, we focus on the problem of robustly building a 3D representation of multi-object scenes from a single RGBD camera image.
Figure 1: Poor understanding of geometry can lead to unwanted collisions during robotic grasping (top), whereas robust, accurate reconstructions can enable precise collision avoidance during manipulation (bottom).
@@
Deep learning methods for 3D reconstruction generally lack a sense of uncertainty about the shape of the object. This can be particularly detrimental when an object in the scene is only partially observed and we seek to factor in the geometry of observed and occluded regions. Such uncertainty can enable safer and more robust operation in a range of downstream tasks, such as robot grasping [chen2024springgrasp, de2024task], safe motion generation and active learning.
@@
Another common approach to building full 3D representations from a partial view is to build a representation solely from the observed data without considering prior information from mesh datasets. A common example of this is the Gaussian Process Implicit Surface [dragiev2011gaussian] model. A more recent example is V-PRISM [wright2024vprism], which probabilistically maps tabletop scenes without using prior information. These reconstruction methods are more robust to unknown objects because they do not rely on any training distribution. However, this also means that they cannot reconstruct the unobserved backside of objects. Humans have the remarkable ability to infer the geometry of scenes based on prior experience, we seek to imbue robots with the same capability.
@@
In this work, we introduce a novel Bayesian approach for robustly reconstructing multi-object tabletop scenes by leveraging object-level shape priors. We present Bayesian Reconstruction with Retrieval-augmented Priors (BRRP). BRRP is resilient to many of the pitfalls of learning-based methods while still being able to leverage an informative prior to more accurately reconstruct known objects. To further improve efficiency, we introduce the idea of a retrieval-augmented prior, where we retrieve relevant components of our prior distribution based on classification results. We begin with an observed RGBD image with corresponding instance segmentations. Then, we compute an identification result that predicts which objects from our database should be retrieved to use as a prior during reconstruction. We use this prior along with a sampled likelihood to infer a posterior distribution over object shapes. Because we solve for a distribution, we can recover principled uncertainty about each object’s shape. In practice, we use a pre-existing foundation model to perform the identification and use registration to make our prior pose invariant. Figure 2 shows an example scene reconstruction, where one of the objects (the orange juice) is decidedly out-of-distribution and our method is still able to reconstruct it according to the observed point cloud.
Figure 2: BRRP (a), (b) takes an input segmented RGBD image, and (c) retrieves objects to use as a prior to (d) reconstruct the scene as well as (e) capture principled uncertainty about object shape, in this case, an out-of-distribution orange juice carton.
@@
We conduct experiments on BRRP in both procedurally generated scenes and in the real world. We quantitatively show that BRRP results in accurate reconstructions and that BRRP is robust to unknown objects in the generated scenes. We qualitatively show that BRRP is robust to noisy real world scenes collected from an RGBD camera, and is able to capture uncertainty within the reconstructions. We also demonstrate downstream improvement with real-world grasping experiments, where the robot is able to more robustly arrive at an accurate pre-grasp pose.
Our three primary contributions can be summarized as:
@@
We organize our paper as follows. We overview related works in Section II. In Section III, we cover mathematical preliminaries for our method. Section IV introduces retrieval-augmented priors. BRRP is introduced in Section V. Experiments are found in Section VI and a conclusion in Section VII
@@ II Related Works
3D Representations. There are many different ways of representing 3D geometry of a scene. In the mapping literature, memory-intensive voxels are used as a representation of the environment. Hilbert maps [ramos2016hilbert], on the other hand, are a continuous occupancy map that takes the form of a linear function over some hinge point feature space. Hilbert map representations have also been extended to Bayesian Hilbert maps of various forms [senanayake2017bayesian, senanayake2018automorphing, zhi2019continuous]. Neural implicit functions have also been used to represent continuous 3D geometry [park2019deepsdf, mescheder2019occupancy, sitzmann2020implicit]. Other representations are built using differentiable rendering and combining multiple views [mildenhall2021nerf, kerbl20233d]. Foundation models have also been developed for this task in [Wang_2024_CVPR] and applied to robotics [sim_geo_pose2024]. We instead focus on the harder problem of reconstruction from only a single-view.
@@
3D Reconstruction with Deep Learning Priors. Many methods have been proposed as ways to leverage deep learning to reconstruct scenes or objects. While some methods aim to predict object shape from RGB data only [xu2024instantmesh, liu2024one, engelmann2021points], we instead focus on using depth measurements during reconstruction. DeepSDF [park2019deepsdf] is a method to reconstruct an object by running inference-time optimization to recover a latent code for a neural implicit function. In the context of robotics, [liao2024uncertainty] extends DeepSDF to have uncertainty-awareness. Other work, such as occupancy networks [mescheder2019occupancy] or PointSDF [van2020learning] try to directly predict such a latent code without inference-time optimization. Deep learning has also been leveraged to learn kernels, which are used to construct a continuous signed distance function [williams2022neural, huang2023neural]. Language is also used during reconstruction in [kasten2024point] and [cheng2023sdfusion]. Another work uses a voxel-to-voxel variational autoencoder conditioned on bounding boxes [saund2021diverse]. While these works typically focus on single object scenes, other work focuses on reconstructing multi-object scenes. For example, [agnew2021amodal] learns to reconstruct a voxel representation, leveraging different input channels to account for occlusion. Alternatively, work such as [iwase2024zero] aim to reconstruct the scene-wide geometry without bothering to separate each object. Instead, we focus on reconstructing individual objects in a manner robust to occlusion. In practice, these deep learning approaches can struggle to reconstruct noisy scenes with multiple, highly occluded, unknown objects on real world depth cameras. These methods can also struggle when segmentation is inaccurate.
@@
3D Reconstruction without Deep Learning Priors. There are also many approaches to perform probabilistic 3D reconstruction without deep learning. Some of such methods for reconstruction use informative prior information by assuming fixed classes of objects, such as 3DP3 [gothoskar20213dp3]. Other methods use an uninformative prior, such as Gaussian process implicit surfaces (GPIS) [dragiev2011gaussian]. While there is an extension of GPIS to a slightly more informative prior [martens2016geometric], the only priors that can be enforced are specifically spherical, ellipsoidal, cylidrical, or planar priors. We instead derive our prior from pre-existing mesh data. V-PRISM [wright2024vprism] is another method that probabilistically maps the scene using a multi-class framing.
@@
Using Reconstructions in Manipulation. 3D reconstruction methods have seen extensive use in manipulation. In [van2020learning], PointSDF provides collision constraints during grasping. PointSDF is also utilized in [matak2022planning], where tactile sensors are used along with the reconstruction during grasping. A learning-based voxel representation is used for grasping in [lundell2019robust]. Neural shape completion is also used during the anthropomorphic grasping pipeline proposed in [hidalgo2023anthropomorphic]. GPIS is also a common representation for manipulation applications. Some recent work has utilized the uncertainty from GPIS representations during grasp selection [chen2024springgrasp, de2024task]. We believe BRRP provides principled uncertainty measurements that can similarly be utilized in downstream manipulation tasks.
@@ III-A Hilbert Maps
A Hilbert map [ramos2016hilbert] is a continuous occupancy map. It represents the environment by a continuous function that is defined by a linear function of a fixed feature transform. Typically this feature transform is induced by a set of hinge points, $\{\mathbf{h}_{1},...,\mathbf{h}_{H}\}\subset\mathbb{R}^{3}$ and a translation-invariant kernel $k(d)$. The transform is then defined as:
@@
Figure 3: (a) A hinge point feature transform induced by a set of hinge points is used by Hilbert maps [ramos2016hilbert]; (b) these maps are built by first sampling negative samples along the unoccupied portions of the camera ray.
@@
Figure 3 shows an illustration of both the hinge point feature transform and the sampling used for creating Hilbert maps.
Hilbert maps have previously been extended to Bayesian Hilbert maps, where a distribution over weights is modeled as a multivariate Gaussian [senanayake2017bayesian, senanayake2018automorphing, zhi2019continuous]. There is also a multiclass variant that defines a weight matrix with Gaussian rows described in [wright2024vprism]. In this work, we adopt the Hilbert map representation, but model the distribution over weights as a collection of particles. This allows our method the capability to capture irregular, non-Gaussian posterior distributions.
@@ III-B Stein Variational Gradient Descent
Figure 4: Left: Overview of BRRP method. We begin with a segmented RGBD image and compute both a retrieval-augmented prior (Section V-A and Section V-B) as well as a likelihood from negative sampling (Section V-C). Then, we combine both in a Bayesian manner via SVGD (Section V-D). Right: An overview of our retrieval-augmented priors for reconstruction. We use CLIP to retrieve the top-$k$ relevant objects in our database (Section V-B), then we perform registration and retrieve the corresponding prior samples (Section V-A).
Stein Variational Gradient Descent (SVGD) [liu2016stein] is an algorithm for variational inference that closely resembles gradient descent. The general problem of variational inference is to find a distribution $q^{*}\in\mathcal{Q}$ that is close to some target distribution $p$. Usually, this takes the form of an optimization problem over the Kullback-Leibler (KL) divergence:
@@
SVGD aims to iteratively transform $q$ in descent directions of the KL divergence in a $d$-dimensional reproducing kernel Hilbert space, $\mathcal{H}^{d}$. Because this Hilbert space is a space of functions, a descent direction requires deriving the functional gradient of our KL divergence objective.
@@
From [liu2016stein]. Let $T(\mathbf{x})=\mathbf{x}+f(\mathbf{x})$, where $f\in\mathcal{H}^{d}$ and $q_{[T]}$ is the density of random variable $\mathbf{z}=T(\mathbf{x})$ when $\mathbf{x}\sim q$. Then
@@
The result of these iterations is that the set of particles converges to an approximation of the target distribution $p$. Importantly, Equation 2 only relies on the gradient of the log of $p$, which means we can perform variational inference to an unnormalized distribution. Such unnormalized distributions are common in many Bayesian inference problems.
@@ IV Retrieval-Augmented Priors
Retrieval-augmented generation [lewis2020retrieval] was originally introduced in the context of improving language generation. The work has served as inspiration for an approach to affordance-prediction in [kuang2024ram]. In our case, we draw inspiration from retrieval-augmented generation, but we use the retrieved results to improve efficiency in certain explicit formulations for prior density functions during Bayesian inference.
To motivate retrieval-augmented priors, consider the problem of Bayesian inference with a mixture model acting as the prior distribution. Given some data, we would like to infer a posterior distribution over hypotheses. Defining a mixture model prior over components $C=\{c_{1},\ldots,c_{N_{c}}\}$ gives:
$P(H|D)\propto P(D|H)\textstyle{\sum_{c\in C}P(H|c)}.$ (3)
If our prior distribution has a lot of components, it may be inefficient to fully evaluate. This could be a serious problem for algorithms like SVGD, which requires iteratively computing the gradient of both the likelihood and prior. Inspired by [lewis2020retrieval], the insight behind retrieval-augmented priors is to determine which subset of the prior distribution components to retrieve and use given some detection result $R$. Conditioning on this detection result, we have a new posterior distribution, $P(H|D,R)$. Making an independence assumption,
@@
$P(H|D,R)\propto P(D|H)\textstyle{\sum_{c\in\text{topk}}P(H|c)P(c|R)}$ (4)
Thus, we only evaluate a subset of the prior components.